OpenStudy (anonymous):

Sqrt -8x+6-3=8x... I need help solving !

OpenStudy (oaktree):

Do you mean\[\sqrt{-8x+6} -3 = 8x\]

OpenStudy (anonymous):

Yes , I need all possible solutions .

OpenStudy (oaktree):

@Gabriel.Dalbani, stop just plugging things into WolframAlpha.

OpenStudy (oaktree):

So what you need to do to solve, @Shgre17, is first move the -3 over to the right-hand side of the equation and then square both sides. So it should look like this:\[-8x + 6 = (8x + 3)^2\]

OpenStudy (oaktree):

Following so far?

OpenStudy (oaktree):

And thanks for the medal, @CGGURUMANJUNATH. I appreciate it.

OpenStudy (anonymous):

Yes

OpenStudy (cggurumanjunath):

you're welcome @OakTree

OpenStudy (oaktree):

Great. So then we FOIL out the right hand side. Can you do that?

OpenStudy (anonymous):

Yes I can, give me a sec to do the math @oaktree

OpenStudy (cggurumanjunath):

wolfram alfa plugging directly gives wrong answer interpreting would be difficult.

OpenStudy (anonymous):

After you foil you would end up with 16x^2+24x+24x+9 right?

OpenStudy (oaktree):

Not quite. Your first term is wrong.

OpenStudy (oaktree):

You see your mistake?

OpenStudy (anonymous):

Lol sorry its 64x^2

OpenStudy (oaktree):

Yeah. Now add the two middle terms and you get...?

OpenStudy (anonymous):

48x

OpenStudy (oaktree):

Right. So we have\[-8x + 6 = 64x^2 + 48x + 9\]Now combine like terms so we have a quadratic equation. We get\[64x^2 + 56x + 3 = 0\]Which doesn't factor nicely, so we have to use the quadratic formula. Do you know it?

OpenStudy (anonymous):

It's ax^2+bx+c right ?

OpenStudy (oaktree):

Not quite. The quadratic formula is \[x = \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

Okay , so what do I do next?

OpenStudy (oaktree):

Well, you take the values for a, b, and c from your equation and plug them in. The two answers you get (since it's plus and minus) are your roots. Okay?

OpenStudy (oaktree):

So it would go like this:\[x = \frac{ -56 \pm \sqrt{56^2 - 4(64)(3)} }{ 2(3) }\]

OpenStudy (anonymous):

Okay I have that

OpenStudy (oaktree):

Which, simplified, is \[\frac{ -56 \pm 8 \sqrt{37} }{ 6 }\]

OpenStudy (oaktree):

Now, do you think you can simplify that all the way and get your answers?

OpenStudy (anonymous):

How did you get that ?

OpenStudy (oaktree):

Simplify the inside of the square root.